### v1

@@ III-C Jaccard score
A score often used in the evaluation of generative models is the Jaccard score or intersection-over-union (IoU) [32, 26]. It has the desirable range of 0 to 1. Of note, the Jaccard score does not exhibit the outsize sensitivity to outliers that the earth-mover’s score has, since the relative distance between non-intersecting densities does not enter into the equation (Methods). In our analysis, the Jaccard score was found to behave like the correlation score in that values were similar regardless of whether fits were high- or low-quality (Table I). The main differences were (1) the range of scores was lower than for the correlation score and (2) one of the low-quality fits—Fig. 2d, with a Jaccard score of 0.644—scored  0.2 lower than the others, which ranged from 0.811 to 0.870. Importantly, the score of 0.834 for the high-quality fit (Fig. 2e) was actually lower than for one of the low-quality fits (0.870 for Fig. 2a), representing grade inflation for the low-quality fit. We found that the Jaccard score is especially prone to grade inflation when one distribution completely circumscribes the other. Together, these results support the conclusion that, like the correlation and earth-mover’s scores, the Jaccard score is not a dependable discriminant of quality in generative models, due to grade inflation.
@@ III-D Kullback-Leibler score
A third relevant quality score is based on the Kullback-Leibler (KL) divergence, also known as relative entropy. This quantity is interpreted as the information lost, or the “surprise,” in approximating one distribution by another [33, 34]. It is used widely in generative models, although in our reading of the literature not as frequently for evaluating synthetic data as the correlation, earth-mover’s, or Jaccard scores. However, the utility of the KL divergence comparing two-dimensional distributions makes it a natural comparator.
@@
While the KL scores from Table I are not that high in absolute terms, the fit for Fig. 2a suffers grade inflation relative to the fits for Figs. 2c-d, which are better fits by eye but have similar scores. The reason for the grade inflation in this case can be traced to the real and synthetic data having similar support but different numbers of modes [34], as follows. Recall that KL divergence can be understood as the amount of surprise at the real data, given the synthetic data. In Fig. 2a, the synthetic data has a single mode, with a support that spans the the real data’s two main modes. Even though the fit is poor, it does have density that overlaps the centers of the two main modes in the real data. As a result, the surprise factor at finding density in these regions of the real data is low, inflating the KL score (which varies inversely with the divergence), despite the low-quality fit. This inconsistency is a drawback of the KL score.
@@ III-E Eden score
The Eden score differs qualitatively from the correlation, earth-mover’s, Jaccard, and KL scores in being an equidensity score, whence its name (see Discussion). Eden is based on the principle that a distribution $p$ is a good match of a distribution $q$ if their regions of probability density $i$ coincide, for all densities $i$. Any reasonable method can be used for determining coincidence; we use the Jaccard score, on the reasoning that Jaccard’s grade inflation is due primarily to its inability to deal appropriately with distributions’ variable densities, and therefore comparisons of equidensity regions would be unlikely to have this problem. To calculate the Eden score, the Jaccard score is calculated for each annulus or ring of density $i$ and these scores are then averaged (Fig. 3). All contours of the distribution count equally, avoiding the issues described above with the earth-mover’s and Jaccard scores (and, of course, with the correlation score).
Figure 3: The Eden score. The Eden score comparing the blue and yellow distributions in the high-quality fit in (a) is calculated as the mean intersection-over-union for each equidensity contour (ring, annulus) (b). “Peaks,” “slopes,” and “foothills” contribute equally. Top row; intersections; bottom row, unions (both in green). The ratios for the three contour levels shown are 0.77, 0.81, and 0.78 (left-to-right), which average to an Eden score of 0.79 for the fit of the two distributions in (a). For clarity, the score is calculated over three contours, instead of the five used in the rest of this study. c-d: Similar for a low-quality fit (Eden score, 0.02).
Indeed, calculation of the Eden score for each of the fits in Fig. 2 showed that it avoids grade inflation (Table I). Scores for the low-quality fits in Fig. 2a-d ranged from 0.162 to 0.261 vs. 0.853 for the high-quality fit in Fig. 2e. (The confidence intervals in Table I account for the stochastic element in our implementation, which can be made arbitrarily small by scaling up the Monte Carlo. The same is true for the KL score. The others are deterministic.) Eden was the only score to demonstrate a consistent gap between low- and high-quality fits, without exceptions. Moreover, this gap was sizable, at 3-4x. Thus, it was immune to grade inflation in these examples.
@@ III-F Validation
Human visual inspection leverages the visual cortex’s millions of years of evolved expertise at comparing sizes and shapes to provide a gold-standard assessment of the similarity two-dimensional distributions. Quantitative expertise hones this ability. Therefore, to validate our results, we showed 20 scientists, data scientists, and engineers 13 fits from pairs of generative models and asked them to choose the better fit. We then compared their choices to each of the five scores we evaluated: correlation, earth-mover’s, Jaccard, KL, and Eden. Raters were not shown any scores, making this a blinded test.
Figure 4: Validation. 20 human experts were asked to rate which of two fits was better for several examples. These ratings were then compared against the five statistical scoring methods. Agreement between each human rater and the score was measured by Cohen’s $\kappa$. a: Percent of raters who agreed the most with each score (Eden, KL, etc.). Mann-Whitney U p-value shown for ranks. b: All $\kappa$ values for each score; Mann-Whitney U p-values shown (Methods). c: Agreement between scores for test pairs (again measured by $\kappa$).
Eighty percent of raters agreed most closely with the Eden score (Fig. 4a), vs. 15% for KL (MWU $p=1.6\times 10^{-4}$), 5% for Jaccard, and 0% for the earth-mover’s and correlation scores. Although Eden consistently outperformed KL on a person-by-person basis (Fig. 4a), its advantage over KL was narrow overall, with a median Cohen’s $\kappa$ of 0.722 across raters (the “excellent” range; 10th–90th percentile: 0.429–0.852) vs. 0.606 (“moderate;” 0.360–0.727) for KL. The similarity in $\kappa$ was because Eden and KL agreed on most pairs (an artifact of how the pairs were chosen). When the analysis was repeated on only those pairs on which Eden and KL disagreed (not shown), 80% of raters agreed with Eden over KL, consistent with Fig. 4a. Both Eden and KL substantially outperformed the earth mover’s score (median $\kappa=0.311$; “low;” 10th-90th percentile 0.129–0.634), correlation score (median $\kappa=-0.019$; -0.126–0.221), and Jaccard score (median $\kappa=0.055$; -0.067–0.222). Of note, the near-zero median $\kappa$s of the correlation scores and Jaccard scores indicate that agreement between these scoring methods and humans with domain expertise is no better than random chance, arguing against their use. In contrast, Eden’s performance matched human intuition quite well. Together with its avoidance of grade inflation, which includes both its clear segregation between low- and high-quality fits and its scoring consistency, its performance supports its use for two-dimensional distributions.
@@ III-G Confidence intervals
For the most direct, apples-to-apples comparisons to training data, generated synthetic data should have the same number of datapoints. However, synthetic data is often needed precisely because the training dataset is small; in principle, sampling the equivalent small number of synthetic datapoints can result in comparatively large sampling error in quality scores. However, we found that the alternative of oversampling from the model can result in a synthetic data KDE that is qualitatively sharper than that of the real dataset, potentially complicating comparisons (Fig. 5). To illustrate, we trained a Gaussian copula on a target with a striped pattern (Fig. 5a). The normal KDE plot (i.e. without oversampling of the synthetic data) does not pick up the stripes (Fig. 5b), whereas the KDE plot with oversampled synthetic data does (Fig. 5c). This lowers the Eden score (from 0.452 to 0.160), as might be expected from how different the KDE becomes. Of the five scoring methods, only Eden and KL were sensitive enough to fall in response to this difference. The correlation, earth-mover’s, and Jaccard scores actually rose, but were all $\geq$0.941 to begin with, reflecting the grade inflation problem.
Our results support repeat sampling as a middle ground between sampling error and oversampling: i.e., scoring each sample and reporting summary statistics such as the mean ± standard deviation of the scores, the median, interquartile scores, or simply the $n$th percentile score. The latter is more conservative than the mean or median: it is the score that $n$ percent of the scores are better than. Figs. 6a and 6c show the distribution of values for each score for 5,000 resamplings for the low-quality fit from Fig. 2a and the high-quality fit from Fig. 2e, respectively. Notably, the distributions spanned at least several percent for all five scores. The earth-mover’s score on the low-quality fit was especially variable, with an interquartile range of $0.22$. The extent of the variation observed supports caution in grading generative models, and strongly suggests that results from single samples can be misleading, may lack discriminative power, and therefore should be avoided if possible.
@@
Figure 6: Confidence intervals. a: Score distributions from 5,000 repeat samples for the fit from Fig. 2a, a low-quality fit; 5th- and 95th-percentile limits shown. Note, all 5,000 samples are from the same fit. b: Representative KDE plots for samples drawn from different parts of the Eden score distribution (with Eden scores shown). c: Score distributions from 5,000 repeat samples for the fit from Fig. 2e, a high-quality fit, with 5th- and 95th-percentile limits. Again, all 5,000 samples are from the same fit. d: Representative KDE plots for samples drawn from different parts of the Eden score distribution (with Eden scores shown).
@@ IV Discussion
All modeling benefits from reliable quality scores, including two-dimensional distributions resulting from generative models. Scores can mislead in several ways, including overfitting and in selecting the wrong model [35]. Here we add the grade-inflation problem, named for a perenially-decried phenomenon in U.S. higher education [36], in which a statistic scores a model higher than it should. We describe why the commonly used correlation score should be particularly prone to grade inflation, and offer examples from real-world datasets where other commonly used scores—specifically, the earth-mover’s, Jaccard, and KL scores—still have this problem, while at least one other score—our newly proposed Eden score—appears not to. We also show the value of multiple sampling for measuring and reporting confidence in these scores, which appears to be a relatively uncommon practice in the literature. We suspect the grade-inflation problem is not new, but is made newly relevant by the explosive growth of data science, the need to select among high-performing models, the bottleneck human visual inspection presents in the setting of such growth in interest, and the computational resources available for data visualization (for example, pairwise scatter plots for large, high-dimensional datasets, which taxed previous generation of computers, sometimes prohibitively).
It is interesting that the Eden score agreed substantially better with human perception of goodness-of-fit than the other scores tested, in a blinded head-to-head comparison on data from generative models. This finding suggests that when human raters compare distributions for similarity, they, like the Eden score, might also be comparing distributions at multiple densities and subconsciously averaging the result. The caveats: raters were specifically asked to consider all contours, were shown the distributions as KDE plots (as opposed to, for example, as three-dimensional and/or interactive renderings), and were limited to people with scientific backgrounds. This is an interesting topic for further investigation.
@@
What explains the difference between the earth-mover’s, Jaccard, and KL scores, which exhibited grade inflation, and Eden, which did not? One answer is that these two sets of scores differ qualitatively in how they weight different regions of the two distributions. The earth-mover’s score weights each datapoint equally, regardless of where in a distribution that point lies. Because by definition there are more datapoints in areas of higher density, the earth-mover’s score will tend to be high as long as the regions of highest density line up well between the two distributions, almost regardless of how mismatched the low-density regions are (in the absence of extreme outliers). The Jaccard and KL scores also weight each datapoint equally, with similar results. Note, these three scores are likely not the only ones with this property; for example, multi-dimensional extensions of the Kolmogorov-Smirnov [KS] statistic also seem to have this property [37], and the Frechet inception distance is likely to suffer grade inflation as well, since it is based on the earth-mover’s distance (as possibly scores based on other $f$-divergences). We coin the term “equipoint” for scores of this kind. Based on these examples, we hypothesize that all equipoint scores will suffer from this tyranny of the majority, yielding high scores even when low-density regions are poorly fit, resulting in grade inflation (Fig. 2 and Table I).
@@
In contrast, the Eden score weights datapoints unequally, such that all densities contribute equally to the score. In such “equidensity” scores, the region of lowest density carries equal weight to the region of highest density, and indeed to all density regions in between (Fig. 3). Topologically, both the peaks and the foothills of the real and synthetic distributions have to line up well to get a high equidensity score, whereas only the peaks need to line up to get high equipoint scores. We propose that equidensity scores such as Eden should be preferred whenever accuracy is required throughout the distribution. (We exclude “unreasonable” equidensity scores from consideration, e.g. taking a trivially large $n$th root of Eden such that all scores end up arbitrarily close to 1.) For example, for medical applications, it is often just as important that generative models accurately capture the features of rare presentations (foothills) as common ones (peaks) [7]. Since one of the main purposes of generative models is to fill out rare cases, we conclude that equidensity scores will generally have the advantage over equipoint scores for this application. Although to our knowledge the Eden score is the first equidensity score, or at least the first recognized as such, it is possible that scores based on other measures that up-weight tails also have something of an equidensity character (e.g. the one-dimensional Anderson-Darling statistic, which up-weights tails more than the KS and e.g. Cramér-von Mises statistics and can be extended to two dimensions [38]). This is a topic for future study.

### v2

@@ III-C Jaccard score
A score often used in the evaluation of generative models is the Jaccard score or intersection-over-union (IoU) [32, 26]. It has the desirable range of 0 to 1. Of note, the Jaccard score does not exhibit the outsize sensitivity to outliers that the earth-mover’s score has, since the relative distance between non-intersecting densities does not enter into the equation (Methods). In our analysis, the Jaccard score was found to behave like the correlation score in that values were similar regardless of whether fits were high- or low-quality (Table I). The main differences were (1) the range of scores was lower than for the correlation score and (2) one of the low-quality fits—Fig. 2d, with a Jaccard score of 0.644—scored  0.2 lower than the others, which ranged from 0.811 to 0.870. Importantly, the score of 0.834 for the high-quality fit (Fig. 2e) was actually lower than for the lowest-quality fit (0.870 for Fig. 2a), indicating grade inflation. We found that the Jaccard score is especially prone to grade inflation when one distribution completely circumscribes the other. Together, these results support the conclusion that, like the correlation and earth-mover’s scores, the Jaccard score is not a dependable discriminant of quality in generative models, due to grade inflation.
@@ III-D Kullback-Leibler score
A third relevant quality score is based on the Kullback-Leibler (KL) divergence, also known as relative entropy. This quantity is interpreted as the information lost, or the “surprise,” in approximating one distribution by another [33, 34]. It is used widely in generative models, although in our reading of the literature not as frequently for evaluating synthetic data as the correlation, earth-mover’s, or Jaccard scores. However, the ubiquity of the KL divergence and its utility comparing two-dimensional distributions make it a natural comparator.
@@
While the KL scores from Table I are not that high in absolute terms, the fit for Fig. 2a suffers grade inflation relative to the fits for Figs. 2c-d, which are better fits by eye but have similar scores. The reason for the grade inflation in this case is because the real and synthetic data have similar support but different numbers of modes [34], as follows. Recall that KL divergence can be understood as the amount of surprise at the real data, given the synthetic data. In Fig. 2a, the synthetic data has a single mode, with a support that spans the the real data’s two main modes. Even though the fit is low quality, it does have density that overlaps the centers of the two main modes in the real data. As a result, the surprise factor at finding density in these regions of the real data is low, inflating the KL score (which varies inversely with the divergence), despite the low-quality fit. This inconsistency is a drawback of the KL score.
@@ III-E Eden score
The Eden score differs qualitatively from the correlation, earth-mover’s, Jaccard, and KL scores in being an equidensity score, whence its name (see Discussion). Eden is based on the principle that a distribution $p$ is a good match of a distribution $q$ if their regions of probability density $i$ coincide, for all densities $i$. Any reasonable method can be used for determining coincidence; we use the Jaccard score, since Jaccard exhibits grade inflation only when densities vary, which is definitionally not the case for equidensity regions; therefore unlike when comparing complete distributions, comparison of equidensity regions should not have this problem. To calculate the Eden score, the Jaccard score is calculated for each annulus or ring of density $i$ and these scores are then averaged (Fig. 3). All contours of the distribution count equally, avoiding the issues encountered with the earth-mover’s and Jaccard scores (and, of course, with the correlation score).
Figure 3: The Eden score. The Eden score comparing the blue and yellow distributions in the high-quality fit in (a) is calculated as the mean intersection-over-union for each equidensity contour (ring, annulus) (b). Peaks, slopes, and foothills contribute equally. Top row; intersections; bottom row, unions (both in green). The ratios for the three contour levels shown are 0.77, 0.81, and 0.78 (left-to-right), which average to an Eden score of 0.79 for the fit of the two distributions in (a). For clarity, the score is calculated over three contours, instead of the five used in the rest of this study. c-d: Similar for a low-quality fit. The peaks are disjoint (ratio, 0.00), the slopes intersect by only a sliver (ratio, <0.01), and the foothills’ intersection-over-union is 0.09, yielding an Eden score of 0.03.
Indeed, calculation of the Eden score for each of the fits in Fig. 2 showed that it avoids grade inflation (Table I). Scores for the low–quality fits in Fig. 2a-d ranged from 0.162 to 0.261 vs. 0.853 for the high-quality fit in Fig. 2e. (The confidence intervals in Table I account for the stochastic element in our implementation, which can be made arbitrarily small by scaling up the Monte Carlo; the same is true for the KL score. The others are deterministic.) Eden was the only score to demonstrate a consistent gap between low- and high-quality fits, without exceptions. Moreover, this gap was sizable, at 3-4x. Thus, it was immune to grade inflation in these examples.
@@ III-F Validation
Human visual inspection leverages the visual cortex’s millions of years of evolved expertise at comparing sizes and shapes to provide a gold-standard assessment of the similarity of two-dimensional distributions. Quantitative expertise hones this ability. Therefore, to validate our results, we showed 20 scientists, data scientists, and engineers 13 fits from pairs of generative models and asked them to choose the better fit. We then compared their choices to each of the five scores we evaluated: correlation, earth-mover’s, Jaccard, KL, and Eden. Raters were not shown any scores, making this a blind test.
Figure 4: Validation. 20 human experts were each asked to rate which of two fits was better for several pairs. Ratings were then compared against each of the five statistical scoring methods. Agreement between each human rater and the scoring method was measured by Cohen’s $\kappa$. a: Percent of raters who agreed the most with each scoring method (Eden, KL, etc.). p-value is for Mann-Whitney U on the ranks. b: All $\kappa$ values for each score, with Mann-Whitney U p-values (Methods). c: Agreement between methods for test pairs (again measured by $\kappa$).
Eighty percent of raters agreed most closely with the Eden score (Fig. 4a), vs. 15% for KL (MWU $p=1.6\times 10^{-4}$), 5% for Jaccard, and 0% for the earth-mover’s and correlation scores. Although Eden consistently outperformed KL on a person-by-person basis (Fig. 4a), its advantage over KL was small overall (Fig. 4b), with a median Cohen’s $\kappa$ of 0.722 across raters (the “excellent” range; 10th–90th percentile: 0.429–0.852) vs. 0.606 (“moderate” agreement; 0.360–0.727) for KL. The explanation: Eden and KL happened to agree on most pairs of fits in the test (Fig. 4c), an artifact of how the pairs were chosen (Methods). When the analysis in Fig. 4b was repeated on only those pairs on which Eden and KL disagreed (not shown), 80% of raters agreed with Eden over KL, consistent with Fig. 4a.
Both Eden and KL substantially outperformed the earth mover’s score (median $\kappa=0.311$; “low” agreement, 10th-90th percentile 0.129–0.634), correlation score (median $\kappa=-0.019$; -0.126–0.221), and Jaccard score (median $\kappa=0.055$; -0.067–0.222). Of note, the near-zero median $\kappa$s of the correlation scores and Jaccard scores indicate that agreement between these scoring methods and humans with domain expertise is no better than chance, arguing against their use. In contrast, Eden’s showed excellent agreement with human gold standard. Together with its avoidance of grade inflation, including both its clear segregation between low- and high-quality fits and its scoring consistency, Eden’s performance supports its use for evaluating two-dimensional distributions.
@@ III-G Confidence intervals
For the most direct, apples-to-apples comparisons, generated synthetic data should have the same number of datapoints as the training data. However, synthetic data is often needed precisely because the training dataset is small; in principle, sampling the equivalent small number of synthetic datapoints can result in comparatively large sampling error in quality scores. However, we found that the alternative of oversampling from the model can result in a synthetic-data KDE that is qualitatively sharper than that of the real dataset, potentially complicating comparisons (Fig. 5). To illustrate, we trained a Gaussian copula on a target with a striped pattern (Fig. 5a). The normal KDE plot (i.e. without oversampling of the synthetic data) does not pick up the stripes (Fig. 5b), whereas the KDE plot with oversampled synthetic data does (Fig. 5c). This lowers the Eden score (from 0.452 to 0.160), as might be expected from how different the KDE becomes. Of the five scoring methods, only Eden and KL were sensitive enough to fall in response to this difference. The correlation, earth-mover’s, and Jaccard scores actually rose, but were all $\geq$0.941 to begin with, reflecting the grade inflation problem.
Our results support repeat sampling as a middle ground between sampling error and mismatch due to oversampling: i.e., scoring each sample and reporting summary statistics such as the mean ± standard deviation of the scores, the median, interquartile scores, or simply the $n$th percentile score. The latter is more conservative than the mean or median: it is the score that $n$ percent of the scores are better than. Figs. 6a and 6c show the distribution of values for each score for 5,000 resamplings for the very low-quality fit from Fig. 2a and the high-quality fit from Fig. 2e, respectively. (Figs. 6b and 6d show KDEs from different samples drawn from the range of scores for Eden.) Notably, the score distributions spanned at least several percent for all five scoring methods. The earth-mover’s score on the very low-quality fit was especially variable, with an interquartile range of $0.22$ (5th-95th percentiles, $0.23$-$0.72$). The extent of the variation observed in Fig. 6 supports caution in grading generative models, and strongly suggests that results from single samples can be misleading, may lack discriminative power, and therefore should be avoided if possible.
@@
Figure 6: Confidence intervals. a: Score distributions from 5,000 repeat samples for the fit from Fig. 2a, a very low-quality fit; 5th- and 95th-percentile limits shown. Note, all 5,000 samples are from the same fit. b: Representative KDE plots for samples drawn (purple lines) from different parts of the Eden score distribution (with Eden scores shown). c: Score distributions from 5,000 repeat samples for the fit from Fig. 2e, a high-quality fit, with 5th- and 95th-percentile limits. Again, all 5,000 samples are from the same fit. d: Representative KDE plots for samples drawn from different parts of the Eden score distribution (with Eden scores shown).
@@ IV Discussion
All modeling benefits from reliable quality scores, including two-dimensional distributions resulting from generative models. Scores can mislead in several ways, including overfitting and in leading to selection of the wrong model [35]. Here we add to these dangers the grade-inflation problem, named for a perenially-decried phenomenon in U.S. higher education [36], in which a statistic scores a model higher than it should. We describe why the commonly used correlation score should be particularly prone to grade inflation, and offer examples from real-world datasets where other commonly used scores—specifically, the earth-mover’s, Jaccard, and KL scores—still have this problem, while at least one other score, our newly proposed Eden score, appears not to. We also show the value of multiple sampling for measuring and reporting confidence in these scores, which appears to be a relatively uncommon practice in the literature. We suspect the grade-inflation problem is not new, but is made newly relevant by the explosive growth of data science, the need to select among high-performing models, the inability to keep up via human visual inspection, and the computational resources increasingly available for data visualization (for example, enabling pairwise scatter plots for ever higher-dimensional datasets).
It is interesting that the Eden score agrees substantially better with human perception of goodness-of-fit than the other scores tested, in a blind head-to-head comparison on data from generative models. This finding suggests that when human raters compare distributions for similarity, they, like the Eden score, might also be comparing distributions at multiple densities and subconsciously averaging the result. The caveats: raters were specifically asked to consider all contours, were shown the distributions as KDE plots (as opposed to, for example, as three-dimensional and/or interactive renderings), and were limited to people with scientific backgrounds. This is an interesting topic for further investigation.
@@
What explains the difference between the earth-mover’s, Jaccard, and KL scores on the one hand, which exhibited grade inflation, and Eden, which did not? One answer is that these two sets of scores differ qualitatively in how they weight different regions of the two distributions being compared. The earth-mover’s score weights each datapoint equally, regardless of where in a distribution that point lies. Because by definition there are more datapoints in areas of higher density, the earth-mover’s score will tend to be high as long as the regions of highest density line up well between the two distributions, almost regardless of how mismatched the low-density regions are (absent extreme outliers). The Jaccard and KL scores also weight each datapoint equally, with similar results. Note, these three scores are likely not the only ones with this property. For example, multi-dimensional extensions of the Kolmogorov-Smirnov [KS] statistic also seem to have this issue [37], and the Frechet inception distance is likely to suffer grade inflation as well, since it is based on the earth-mover’s distance. We see no obvious reason that scoring methods based on other $f$-divergences would not also have this problem. We coin the term “equipoint” for scores of this kind. Based on these examples, we hypothesize that all equipoint scores will suffer from this tyranny of the majority, yielding high scores even when low-density regions are poorly fit, resulting in grade inflation (Fig. 2 and Table I).
@@
In contrast, the Eden score weights datapoints unequally, such that all densities contribute equally to the score. In such “equidensity” scores, the region of lowest density carries equal weight to the region of highest density, and indeed to all density regions in between (Fig. 3). Topologically, both the peaks and the foothills of the real and synthetic distributions have to line up well to get a high equidensity score, whereas only the peaks need to line up to get high equipoint scores. We propose that equidensity scores such as Eden should be preferred whenever accuracy is required throughout the distribution. (We exclude “unreasonable” equidensity scores from consideration, e.g. taking a trivially large $n$th root of Eden such that all scores end up arbitrarily close to 1.) For example, for medical applications, it is often just as important that generative models accurately capture the features of rare presentations (foothills) as common ones (peaks) [7]. Since one of the main purposes of generative models is to fill out rare cases, we conclude that equidensity scores will generally have the advantage over equipoint scores for this application. Although to our knowledge the Eden score is the first equidensity score, or at least the first recognized as such, it is possible that scores based on other measures that up-weight tails also have something of an equidensity character; potential examples include the one-dimensional Anderson-Darling statistic, which up-weights tails more than the KS and e.g. Cramér-von Mises statistics, and which can be extended to two dimensions [38]. This is a topic for future study.